Can you please give me a step by step on some of these? 23. Use the substitution method to evaluate the following indefinite integrals. (a) / 22(22 + 1)23 dx : + 7)(x2 + 7x + 3)4/5
Answer #1

a) Jern &3 dx= - Ict u=8+1 du=arde 1. =/ u23 dua ett kézifu+c b) ] 22+77(4772+3) dr = let uz 474+3 du=(2x477 da guls dua Z 9